A 1.40 kg squirrel runs east across the ground for 7.00 meters. She climbs 0.650 meters before running west on a limb for 4.50 meters. She jumps up 0.990 meters onto another branch, travels west 4.55 meters. She stops and eats a nut. Calculate net work. 



1
Expert's answer
2020-12-29T15:29:18-0500

The total work done against gravity is equal to


"W = mgh"

where "m = 1.4kg" is the mass of the squirrel, "g = 9.8m\/s^2" is the gravitational acceleration, and "h" is the total heigth change. From the problem statement we can find, that


"h = 0.650m + 0.990m = 1.64m"

Thus, the net work is


"W =1.4kg\cdot 9.8m/s^2\cdot 1.64m \approx 22.5J"

Answer. 22.5 J